# Hierarchical Long Video Understanding with Audiovisual Entity Cohesion and Agentic Search

### v1

@@ 3.2.4 Scene Segmentation and Global Summary
Long videos often consist of multiple temporally extended scenes that share coherent narratives, recurring characters, and consistent environments. Existing approaches typically segment videos into very short clips (e.g., 5 seconds in [31]) or aggregate fixed-size chunks [19], which may capture isolated moments but fail to model long-range temporal dependencies. To address this, we perform adaptive scene-level aggregation based on the segment-level descriptions $\{C_{i}^{t}\}_{i=1}^{N}$, where $N$ is the total number of segments. These captions are split into overlapping chunks and processed by an LLM to perform semantic grouping, merging consecutive segments into coherent scenes:
@@
where each $s_{j}$ represents a temporally contiguous scene with a consistent narrative focus, and $c_{i}$ denotes the $i$-th segment. The boundaries $(a_{j},b_{j})$ are adaptively determined by the LLM based on the continuity of events observed across captions.
Each resulting scene $s_{j}$ is then summarized by the LLM to produce a concise scene-level description $\tilde{s}_{j}$, capturing its key characters, events, and transitions. The collection of these summaries forms the high-level scene set $\tilde{\mathcal{S}}=\{\tilde{s}_{1},\tilde{s}_{2},\ldots,\tilde{s}_{M}\}$, which provides a structured narrative view of the video. Finally, a global summary $\tilde{\mathcal{G}}$ is generated from $\tilde{\mathcal{S}}$, describing the main storyline, recurring entities, and overall context (e.g., background and video type).
@@ 3.3 Agentic Search with Reasoning
Similar to [31], our agentic search framework adopts an iterative think–act–observe loop for autonomous and adaptive information retrieval. At the core, we use a reasoning LLM as the planner. The reasoning process is tool-augmented: the agent first plans actions and formulates queries (think), then invokes specialized tools with appropriate parameters (act), and finally integrates tool outputs into subsequent reasoning (observe). This synergy between LLM-driven reasoning and tool interaction enables progressive query refinement, evidence-guided navigation, and accurate answer generation.
To provide holistic context, the global summary $\tilde{\mathcal{G}}$ is prepended to each initial prompt, while tool outputs dynamically enrich the agent’s reasoning across iterations. This design highlights the critical role of tools in extending the agent’s capabilities and underscores the importance of tight LLM–tool collaboration for robust and adaptive retrieval.
@@ 3.3.1 Multi-Granularity Tools
We denote the set of all available tools as
@@
Each tool $T_{i}\in\mathcal{T}$ is a callable function that maps a textual query (or question) and optional time ranges to a textual response and corresponding timestamps $T_{i}(q,\tau;D)\rightarrow(r,\tau^{\prime})$, where $q$ denotes the user query, $\tau=[t_{s},t_{e}]$ the optional input time window, $D$ the corresponding database, $r$ the retrieved or generated response, and $\tau^{\prime}$ the associated timestamps.
Global Scene Browse.
The Global Scene Browse tool $T_{\text{scene}}$ supports coarse-grained navigation and scene localization along the video timeline. Given a user query $q$ and the scene collection $D=\tilde{\mathcal{S}}$, it identifies and summarizes the most relevant scenes with an LLM, returning their storyline and corresponding timestamps $\tau^{\prime}$. The agent tends to invoke this tool for complex or ambiguous queries involving multiple events or temporal dependencies.
Segment Caption Search.
The Segment Caption Search tool $T_{\text{caption}}$ performs fine-grained text-based retrieval within specified temporal ranges. Given the user query $q$ and the segment database $D=\tilde{\mathcal{C}}$, the tool retrieves the most semantically relevant segment descriptions $r$ along with their associated time spans $\tau^{\prime}$. This is achieved through cosine similarity matching between the query embedding and pre-computed caption embeddings for all video segments, ensuring efficient and accurate retrieval of localized content.
Segment Visual Search.
To capture visual cues that may be overlooked in textual descriptions, the Segment Visual Search tool $T_{\text{visual}}$ serves as a complementary mechanism to Segment Caption Search. While the latter relies on text embeddings, $T_{\text{visual}}$ leverages cross-modal embeddings generated by the UNITE framework [11], as previously described. This enables retrieval based on rich visual semantics aligned with the query, ensuring that visually salient details are incorporated into the search process.
Entity Search.
The Entity Search tool $T_{\text{entity}}$ supports high-level, entity-centric retrieval across large temporal ranges. Given an entity-related query $q$, the tool first retrieves the top-$K_{1}$ most relevant entities from database $\tilde{\mathcal{E}_{g}}$ based on their descriptions in the pre-computed embedding space. For all segments linked to these entities, it then performs a second-stage reranking to select the top-$K_{2}$ most relevant segments from the entity-centric database $\tilde{\mathcal{E}_{e}}$, using the same query. Finally, the tool applies entity-aware re-captioning to the retrieved segments, generating a coherent response $r$ enriched with precise timestamps $\tau^{\prime}$.
Inspection Tool.
The Inspection tool $T_{\text{inspect}}$ provides fine-grained temporal inspection to support detailed reasoning. It consists of two complementary modules: Clip Caption Inspect ($T_{\text{inspect}}^{\text{tex}}$) and Visual Inspect ($T_{\text{inspect}}^{\text{vis}}$). The $T_{\text{inspect}}^{\text{tex}}$ examines coarse textual descriptions to determine what occurred during a specified time span. For example, for a query such as “What does the protagonist do after he jumps down the stairs?”, the tool inspects subsequent time ranges to identify the protagonist’s next actions after locating the event “he jumps down the stairs” in previous iteration. The $T_{\text{inspect}}^{\text{vis}}$ tool, leverages a VLM to perform precise visual verification within a given time range. Due to frame limitations in the VLM we used, this inspection focuses on short intervals, ensuring accurate visual grounding for fine-grained queries.
In both segment search and inspection tools, we prioritize low-cost text-based retrieval over high-cost visual reasoning with VLMs, enabling our system to achieve efficient and coherent long-term video understanding. Together, these multi-granularity tools effectively support diverse query types across our hierarchical database, enabling flexible and accurate information access.
@@ 3.3.2 Multi-Step Reasoning
Besides $(\mathcal{T},D)$, the agent maintains a context memory $\mathcal{M}$ for the reasoning process, initiated to include the global summary $\tilde{\mathcal{G}}$, i.e. $\mathcal{M}_{0}=\{\tilde{\mathcal{G}}\}$. Let $q_{org}$ denotes the original question. At each reasoning step $t$, the planner dynamically selects to call a tool $T_{a_{t}}$ with a modified query $q_{t}$ with optional time ranges based on the current context: $a_{t}=\pi_{\text{LLM}}(q_{org},\mathcal{T},\mathcal{M}_{t})$, where $\pi_{\text{LLM}}$ denotes the LLM-driven policy that determines which tool to invoke. The tool results $(r_{t},\tau^{\prime}_{t})=T_{a_{t}}(q_{t};D)$ and the tool call context are integrated into the context memory to guide subsequent reasoning within the think–act–observe loop, i.e. $\mathcal{M}_{t}=\{\mathcal{M}_{t-1};(T_{a_{t}},q_{t});(r_{t},\tau^{\prime}_{t})\}$. The iteration continues until it finds the answer or the maximum iteration step is reached. Figure 2 shows an example for a counting-like query that needs long contexts. The agent first calls the Global Scene Browse tool to locate the “second-to-last song” and obtain its content and time span. It then calls Visual Inspect tool to analyze the segment of the song “One" (55:00–57:00), which returns “two singers playing guitar and one playing percussion", enabling the agent to derive the final answer.
4 Experimental Results
Table 1: Comparison results on LVBench. Accuray (%) is reported. Bold indicates the best performance, and underscore denotes the second-best.
@@ 4 Experimental Results
Table 2: Comparison results on various long video benchmarks. We use 2 fps in captioning for LVBench and 0.67 fps for other benchmarks.
Methods LVBench Video-MME (long) LongVideoBench EgoSchema
Overall w/o subs w subs Long-val Val Val
Proprietary VLMs
GPT-4o 48.9 65.3 72.1 60.9 66.7 70.4
OpenAI o3 57.1 64.7 - 60.6 67.5 63.2
Open-source VLMs
Qwen2.5-VL-72B 47.7 63.9 - - 60.7 -
AdaReTake 53.3 65.0 76.4 - 67.0 -
Seed1.5-VL-Thinking-200B 64.6 - - - 74.0 -
RAG and agents
VideoRAG 49.2 - - - - -
VideoTree 28.8 - - - - 67.0
VideoAgent 29.3 - - - - 63.2
VideoLucy 58.8 66.8 - - - -
DVD 74.2 67.3 - 68.6 71.6 76.6
DVD w. subtitle 76.0 - - - - -
Ours 84.1 - 85.5 81.0 83.0 81.6
@@ 4.1 Benchmark and Implementation Details
We evaluate our approach on four widely-used long video understanding benchmarks, including LVBench [23], Video-MME [6], LongVideoBench [28] and EgoSchema [16]. The LVBench, a standard benchmark for long-term video question answering, contains 1,549 questions across 103 videos, with an average duration of 4,101 seconds. The dataset covers six diverse categories: temporal grounding (TG), summarization (Sum), reasoning (Rea), entity recognition (ER), event understanding (EU), and key information retrieval (KIR). With high-quality ground-truth annotations, LVBench provides a reliable and challenging benchmark for evaluating comprehensive long video understanding capabilities. For Video-MME, we evaluate it on its long split, which includes 300 videos and 900 questions, with durations ranging from 30 to 60 minutes. For LongVideoBench, we focus on the long subset in the val split, containing 188 videos and 564 questions, with durations between 900 and 3600 seconds. Most videos in this subset lack audio tracks. It is noted that we utilize the official subtitles without speaker identity on LongVideoBench. EgoSchema serves as a diagnostic benchmark for long-form understanding and reasoning. We evaluate on its val split, which contains 500 videos and 500 questions, with each video lasting three minutes and no audio provided. For simplicity, we use audio streams only when the language is English across all datasets.
Implementation Details
@@
For database construction, we employ GPT-4.1 [1] to generate segment-level captions and summarize scenes and entities. Each 30-second segment is captioned using 20 sampled frames (0.67 fps) to ensure efficiency while maintaining coverage. During agentic search, we use OpenAI o3 [17] as the reasoning planner, which aggregates information from tool outputs and produces answers, with a maximum reasoning depth of 10 steps. For query-aware re-captioning during tool calls, we sample 30 frames per segment (1 fps) to provide sufficient visual context. Additionally, we leverage OpenAI o3 as the VLM used in Visual Inspect tool and set its maximum number of input frames to 50. Unless otherwise specified, the same settings are applied across all datasets.
Table 3: Ablation study of various modules on LVBench. All variants use 0.67 fps in captioning.
Methods Speaker Transcript Hierarchical Visual embed Acc.
@@
4.2 Comparison with State-of-the-Art
We compare our framework against several leading approaches for long video understanding, including proprietary large VLMs [1, 17], open-source VLMs [2, 24, 7], RAG-based systems [19] and video agents [26, 25, 32, 31]. We reproduce the results of VideoRAG [19] on LVBench using the official implementation, while other baseline results are taken directly from published reports.
Table 1 summarizes the comparison results on LVBench. Our approach consistently outperforms all baselines in overall accuracy, with particularly strong gains in reasoning (Rea, a challenging category in all baselines) and temporal grounding (TG). These improvements highlight the effectiveness of our multi-granularity information parsing during offline database construction and the multi-level retrieval tools integrated into our agentic search framework. In entity recognition (ER), our method also surpasses DVD, which relies on global subject registries, demonstrating the robustness of our audiovisual entity cohesion strategy. Notably, our system achieves these results using much coarser temporal segmentation (30-second segments with 20 frames for captioning) and fewer reasoning iterations (maximum of 10) compared to DVD’s finer-grained configuration (5-second segments at 2 fps) and deeper reasoning depth (up to 15 steps), underscoring both the efficiency and scalability of our approach.
As our approach employs coarse temporal segmentation and sampling than DVD, our captions inherently contain fewer details. To assess the impact of this difference in textual richness, we compare our scheme with a variant that samples at 2 fps for each 30‑second segment (denoted as Ours (2 fps) in Table 1). Specifically, this variant generates captions for two 15-second sub-segments (each sampled at 30 frames) and concatenates them into a single segment description. We can see that with a larger sampling rate, the proposed scheme performs consistently better in all categories, owing to more detailed textual descriptions by captioning. We find that increasing the sampling rate consistently improves performance across most categories, attributable to the richer textual descriptions generated by denser captioning.
Table 2 further presents the comparison results on all benchmarks. Our method achieves the highest scores on all benchmarks, reaching 84.1 on LVBench, 85.5 on Video-MME (long), 81.0/83.0 on long/val split of LongVideoBench, and 81.6 on EgoSchema, even though most videos in LongVideoBench and EgoSchema lack audio streams. These results underscore the effectiveness of reasoning across the proposed hierarchical structure with semantic coherence.
Figure 3: Comparison of accuracy and efficiency on six categories. “Acc" and “iter" denote accuracy and the average number of reasoning iterations, respectively.
@@ 4.2 Comparison with State-of-the-Art
4.3 Ablation Study
4.3.1 Ablation on Various Modules
We conduct an ablation study on LVBench to evaluate the contributions of several key components: Audio (Speaker and Transcript), Hierarchical, and Visual embed. “Speaker" and “Transcript" denote the use of speaker labels and transcripts in audio, respectively. “Hierarchical" represents the hierarchical organization of video semantics (segment, entity, scene, and global summary) and “Visual embed" corresponds to the visual retrieval capability enabled by the Segment Visual Search tool.
We compare four variants in Table 3. Ours_clip retains only the segment-level search and two inspection tools, emphasizing reasoning via local grounding. Ours_clip_t further disables the Segment Visual Search from Ours_clip, focusing exclusively on textual retrieval at segment level. Ours_visual removes audio transcripts and speaker annotations while keeping all other components, allowing us to assess the effectiveness of purely visual–structural reasoning. Ours_transcript removes speaker identities by the diarization step and constructs the hierarchical database without any speaker-related information.
As shown in Table 3, removing the hierarchical organization in Ours_clip leads to a substantial performance drop compared to Ours, with accuracy dropping from $81.0$ to $72.8$. This underscores the central role of hierarchical indexing in aggregating multi-granular evidence and supporting effective cross‑segment reasoning. The audio modality also contributes significantly to overall performance. In Ours_visual, the accuracy drops to $71.7$ due to the loss of transcripts and speaker information that not only provide complementary context beyond visual cues but also enhance entity consistency. Indeed, transcripts alone provide meaningful semantic enrichment: Ours_transcript improves by 4% over Ours_visual, indicating that textual audio cues facilitate richer reasoning. Furthermore, incorporating speaker identity yields an additional 5.3% accuracy gain when comparing Ours with Ours_transcript. These results emphasize the effectiveness of both transcripts and speaker labels. Although most questions in LVBench focus on visual content, the audio stream nonetheless strengthens entity coherence, thereby improving semantic consistency within the hierarchical structure. Lastly, Ours_clip_t performs slightly worse than Ours_clip, because it lacks visual embedding-based search, which offers additional grounding capabilities. All these results demonstrate that these components contribute uniquely to the system’s overall effectiveness.

### v2

@@ 3.2.4 Scene Segmentation and Global Summary
Long videos often consist of multiple temporally extended scenes with coherent narratives, recurring entities, and consistent environments. Existing approaches typically segment videos into short fixed-length clips (e.g., 5 seconds in [31]) or aggregate fixed-size chunks [19], which can capture local moments but may miss long-range temporal structure. We therefore perform adaptive scene-level aggregation based on segment descriptions $\{C_{i}^{t}\}_{i=1}^{N}$, where $N$ is the total number of segments. We split the sequence into overlapping chunks and use an LLM to group consecutive and semantically related segments into scenes:
@@
where each $s_{j}$ represents a temporally contiguous scene with a consistent narrative focus, and $c_{i}$ denotes the $i$-th segment. The boundaries $(a_{j},b_{j})$ are adaptively determined by the LLM based on the semantic continuity.
Each scene $s_{j}$ is then summarized by an LLM to produce a concise scene-level description $\tilde{s}_{j}$, capturing its key characters, events, and transitions. The collection of these summaries forms the high-level scene set $\tilde{\mathcal{S}}=\{\tilde{s}_{1},\tilde{s}_{2},\ldots,\tilde{s}_{M}\}$. Finally, a global summary $\tilde{\mathcal{G}}$ is generated from $\tilde{\mathcal{S}}$, describing the main storyline, recurring entities, and overall context (e.g., background and video type).
@@ 3.3 Agentic Search with Reasoning
Similar to [31], our agentic retrieval adopts an iterative think-act-observe loop for adaptive information retrieval. At the core, we use a reasoning LLM as the planner, which formulates intermediate queries (think), invokes specialized tools with appropriate parameters (act), and incorporates tool outputs into subsequent reasoning (observe). This synergy between LLM-driven reasoning and tool interaction enables progressive query refinement, evidence-guided navigation, and accurate answer generation. We initialize the agent with the global summary $\tilde{\mathcal{G}}$, and enrich its context with tool outputs across multiple iterations.
@@ 3.3.1 Multi-Granularity Tools
To support adaptive reasoning over long videos, we equip the agent with a suite of multi-granularity retrieval and inspection tools operating on different levels of the hierarchical database $D$. We denote the toolset as
@@
Each tool $T_{i}\in\mathcal{T}$ is a callable function that takes a textual query $q$ with optional time ranges $\tau=[t_{s},t_{e}]$ and returns a textual response $r$ with associated timestamps $\tau^{\prime}$, i.e., $T_{i}(q,\tau;D)\rightarrow(r,\tau^{\prime})$.
The Global Scene Browse tool $T_{\text{scene}}$ performs coarse navigation over scene summaries $\tilde{\mathcal{S}}$ to localize relevant scenes and time ranges. The Segment Caption Search $T_{\text{caption}}$ retrieves fine-grained evidence from segment descriptions in $\tilde{\mathcal{C}}$ via text embedding matching. The Segment Visual Search $T_{\text{visual}}$ complements caption retrieval with visual-semantic search using cross-modal embeddings by UNITE [11] to capture cues missing from text. The Entity Search $T_{\text{entity}}$ conducts entity-centric retrieval over canonical entities $\tilde{\mathcal{E}}_{g}$ and their linked evidence $\tilde{\mathcal{E}}_{e}$ to gather long-range information about specific entities; and the Inspection tool $T_{\text{inspect}}$ provides localized inspection within specified time spans and consists of two complementary modules: Clip Caption Inspect ($T_{\text{inspect}}^{\text{tex}}$), which examines textual descriptions to identify what occurred during the interval, and Visual Inspect ($T_{\text{inspect}}^{\text{vis}}$), which performs VLM-based visual verification for fine-grained confirmation.
During inference, the agent dynamically selects and composes these tools based on the query intent and current context, prioritizing low-cost text retrieval before higher-cost visual inspection. Detailed tool definitions and implementation are provided in Sec. B of the supplementary material.
@@ 3.3.2 Multi-Step Reasoning
In addition to $(\mathcal{T},D)$, the agent maintains a context memory $\mathcal{M}$, initialized with the global summary $\tilde{\mathcal{G}}$, i.e. $\mathcal{M}_{0}=\{\tilde{\mathcal{G}}\}$. Let $q_{org}$ denote the original question. At reasoning step $t$, the planner dynamically selects to call a tool $T_{a_{t}}$ and forms an intermediate query $q_{t}$ with an optional time window based on the current memory: $a_{t}=\pi_{\text{LLM}}(q_{org},\mathcal{T},\mathcal{M}_{t})$, where $\pi_{\text{LLM}}$ denotes the LLM-driven policy that determines which tool to invoke. The tool results $(r_{t},\tau^{\prime}_{t})=T_{a_{t}}(q_{t};D)$, and the tool call context are integrated into the context memory to guide subsequent reasoning, i.e. $\mathcal{M}_{t}=\{\mathcal{M}_{t-1};(T_{a_{t}},q_{t});(r_{t},\tau^{\prime}_{t})\}$. The process repeats until it finds the answer or a maximum number of steps is reached. Figure 2 shows an example for a counting-like query that needs long contexts. The agent first calls the Global Scene Browse tool to locate the “second-to-last song” and obtain its content and time span. It then calls Visual Inspect tool to analyze the segment of the song “One” (55:00-57:00), which returns “two singers playing guitar and one playing percussion”, enabling the agent to derive the final answer.
Table 1: Comparison on LVBench. Accuracy (%) is reported. Bold indicates the best performance, and underscore denotes the second-best.
@@
4 Experimental Results
Table 2: Comparison on other long video benchmarks. We use 0.67 fps for captioning in our approach. “Video-MME (L)” denotes the long split of Video-MME. “LVB” denotes LongVideoBench.
Methods Video-MME (L) LVB EgoSchema
w/o sub w sub Long Val
GPT-4o 65.3 72.1 60.9 70.4
OpenAI o3 64.7 - 60.6 63.2
Qwen2.5-VL-72B 63.9 - - -
AdaReTake 65.0 76.4 - -
VideoTree - - - 67.0
VideoAgent - - - 63.2
VideoLucy 66.8 - - -
DVD 67.3 - 68.6 76.6
Ours - 82.8 78.2 81.6
@@ 4.1 Benchmark and Implementation Details
Benchmark We evaluate our approach on four widely-used long video understanding benchmarks, including LVBench [23], Video-MME [6], LongVideoBench [28] and EgoSchema [16]. The LVBench contains 1,549 questions across 103 videos, with an average duration of 4,101 seconds. The dataset covers six diverse categories: temporal grounding (TG), summarization (Sum), reasoning (Rea), entity recognition (ER), event understanding (EU), and key information retrieval (KIR). With high-quality ground-truth annotations, LVBench provides a reliable and challenging benchmark for evaluating comprehensive long video understanding capabilities. For Video-MME, we evaluate it on its long split, which includes 300 videos and 900 questions, with durations ranging from 30 to 60 minutes. For LongVideoBench, we focus on the long subset in the val split, containing 188 videos and 564 questions, with durations between 900 and 3600 seconds. Most videos in this subset lack audio tracks and we utilize the official subtitles without speaker identity. EgoSchema serves as a diagnostic benchmark for long-form understanding and reasoning. We evaluate on its val split, which contains 500 videos and 500 questions, with each video lasting three minutes and no audio provided. For simplicity, we use audio streams only when the language is English across all datasets.
@@
Implementation Details For database construction, we employ GPT-4.1 [1] to generate segment-level captions and summarize scenes and entities. Each 30-second segment is captioned using 20 sampled frames (0.67 fps) to ensure efficiency while maintaining coverage. During agentic search, we use OpenAI o3 [17] as the reasoning planner, which aggregates information from tool outputs and produces answers, with a maximum reasoning depth of 10 steps. For query-aware re-captioning during tool calls, we sample 30 frames per segment (1 fps) to provide sufficient visual context. Additionally, we leverage OpenAI o3 as the VLM used in Visual Inspect tool and set its maximum number of input frames to 50. Unless otherwise specified, the same settings are applied across all datasets.
4.2 Comparison with State-of-the-Art
We compare our framework against several leading approaches for long video understanding, including proprietary large VLMs [1, 17], open-source VLMs [2, 24, 7], RAG-based systems [19], and video agents [26, 25, 31, 32]. We reproduce the results of VideoRAG [19] on LVBench using the official implementation, while other baseline results are taken directly from published reports.
Table 1 summarizes the results on LVBench. Our approach consistently outperforms all baselines in overall accuracy, with especially strong gains in reasoning (Rea), a challenging category for existing methods, and temporal grounding (TG). These improvements highlight the effectiveness of our semantically consistent hierarchy coupled with multi-level agentic search. In entity recognition (ER), our method also surpasses DVD, which relies on global subject registries, demonstrating the robustness of our audiovisual entity cohesion strategy. Notably, our system achieves these results using much coarser temporal segmentation (30-second segments with 20 frames for captioning) and fewer reasoning iterations (up to 10) compared to DVD’s finer-grained setup (5-second segments at 2 fps) and deeper reasoning depth (up to 15 steps), underscoring both the efficiency and scalability of our approach. Further increasing the sampling rate to 2 fps for each 30-second segment (denoted as Ours (2 fps) in Table 1) yields an additional 3.1% accuracy improvement, attributed to the richer textual descriptions enabled by denser captioning.
Table 2 further presents the comparison results on other benchmarks. Our method achieves the highest scores on all benchmarks, reaching 82.8% on Video-MME (long), 78.2% on long val split of LongVideoBench, and 81.6% on EgoSchema, even though most videos in LongVideoBench and EgoSchema lack audio streams. These results underscore the effectiveness of reasoning across the proposed hierarchical structure with semantic coherence.
4.3 Ablation Study
4.3.1 Ablation on Various Modules
We conduct an ablation study on LVBench to evaluate the contributions of several key components: the Speaker clues and Transcript in audio stream, the Hierarchical organization of video semantics, and the Visual embed for visual retrieval. We compare four variants in Table 3. Ours_clip retains only the segment-level search and two inspection tools, emphasizing reasoning via local grounding. Ours_clip_t further disables the Segment Visual Search from Ours_clip, focusing exclusively on textual retrieval at segment level. Ours_visual removes audio transcripts and speaker annotations while keeping all other components, allowing us to assess the effectiveness of purely visual-structural reasoning. Ours_trans removes speaker identities by the diarization step and constructs the hierarchical database without any speaker-related information.
Table 3: Ablation study on LVBench. All variants use 0.67 fps in captioning. “Spk.”, “Trans.”, “Hier.”, and “Vis emb.” denote Speaker, Transcript, Hierarchical, Visual embed, respectively.
Methods Spk. Trans. Hier. Vis emb. Acc.(%)
@@ 4.3.1 Ablation on Various Modules
As shown in Table 3, removing the hierarchical organization in Ours_clip leads to a substantial performance drop compared to Ours, with accuracy dropping from $81.0$ to $72.8$. This underscores the central role of hierarchical indexing in aggregating multi-granular evidence and supporting effective cross‑segment reasoning. The audio modality also contributes significantly to overall performance. In Ours_visual, the accuracy drops to $71.7$ due to the loss of transcripts and speaker information that not only provide complementary context beyond visual cues but also enhance entity consistency. Indeed, transcripts alone provide meaningful semantic enrichment: Ours_trans improves by 4% over Ours_visual, indicating that textual audio cues facilitate richer reasoning. Furthermore, incorporating speaker identity yields an additional 5.3% accuracy gain when comparing Ours with Ours_trans. These results emphasize the effectiveness of both transcripts and speaker labels. Although most questions in LVBench focus on visual content, the audio stream nonetheless strengthens entity coherence, thereby improving semantic consistency within the hierarchical structure. Lastly, Ours_clip_t performs slightly worse than Ours_clip, because it lacks visual embedding-based search, which offers additional grounding capabilities. All these results demonstrate that these components contribute uniquely to the system’s overall effectiveness.